OpenStudy (anonymous):

find dxy y=6x-8/(x-8) u= 6x-8 v=x-8 u' = 6 v'=1 (6x-8)1 + (x-8)6 /(x-8)^2 and I get -56 / (x-8)^2 but the answer should be -40 / (x-8)^2 . So what did I do wrong?

OpenStudy (phi):

isn't the rule \[ \frac{d}{dx}\frac{u}{v} = \frac{v\ du - u\ dv}{v^2} \]

OpenStudy (phi):

you are adding , not subtracting

OpenStudy (anonymous):

oh right... but when I do that I get 40 instead of -40 for the numerator.

OpenStudy (phi):

yes, it is v du - u dv (v is the denominator and u is the numerator)

OpenStudy (anonymous):

Okay. SO i just got the formula backwards.

OpenStudy (phi):

yes, I can never remember it.

OpenStudy (phi):

if I get confused I write it as (6x-8) (x-8)^-1 and use the product rule. I think the other way is faster, but you have to remember what gets subtracted.

OpenStudy (anonymous):

True. Thanks for the help!